Vortex Company operates a retail store with two departments. Information about those departments follows.

Department A Department B
Sales $ 838,500 $ 451,500
Cost of goods sold 440,000 293,475
Direct expenses:
Salaries 123,000 86,000
Insurance 16,000 11,000
Utilities 21,000 29,500
Depreciation 16,000 13,000
Maintenance 6,400 5,800


The company also incurred the following indirect costs.

Salaries $26,000
Insurance 6,600
Depreciation 14,600
Office expenses 40,000


Indirect costs are allocated as follows: salaries on the basis of sales; insurance and depreciation on the basis of square footage; and office expenses on the basis of number of employees. Additional information about the departments follows.

Department Square footage Number of employees
A 32,200 77
B 13,800 33


Required:
1. Determine the departmental contribution to overhead and the departmental net income for department A and Department B.
2. Should Department B be eliminated?
VORTEX COMPANY Departmental Contribution Statements Department A Department B Sales $ 838,500 $ 451,500 Cost of goods sold 44Should Department B be eliminated? Should Department B be eliminated?
Answer #1

Answer-1)-

VORTEX COMPANY
DEPARTMENTAL CONTRIBUTION STATEMENTS
PARTICULARS DEPARTMENT A DEPARTMENT B
$ $
Sales 838500 451500
Less- Cost of goods sold 440000 293475
Gross Profit 398500 158025
Less- Direct Expenses
Salaries 123000 86000
Insurance 16000 11000
Utilities 21000 29500
Depreciation 16000 13000
Maintenance 6400 5800
Total Direct expenses 182400 145300
Departmental contribution to overhead 216100 12725
Less- Allocated indirect expenses
Salaries (26000*838500)/1290000=16900 (26000*451500)/1290000=9100
Insurance (6600*32200)/46000=4620 (6600*13800)/46000=1980
Depreciation (14600*32200)/46000=10220 (14600*13800)/46000=4380
Office (40000*77)/110 =28000 (40000*33)/110 =12000
Total indirect expenses 59740 27460
Operating Income (loss) 156360 -14735

2)- Department B should not be eliminated because it provides a positive departmental contribution.